OpenStudy (anonymous):

okay no i have a question ;D it's the write the equation for a piece wise fucntion junk, but i don't get this one.....?.....................

OpenStudy (anonymous):

i will upload a picture ...

OpenStudy (anonymous):

\[f(x) = |x+4| = \left\{\begin{array}{rcc} x + 4 & \text{if} & x \geq -4 \\ - x - 4& \text{if} & x < -4 \end{array} \right. \] just showing off

OpenStudy (anonymous):

haha very funny

OpenStudy (anonymous):

sorry, couldn't resist. i will be quiet and wait for the picture

OpenStudy (anonymous):

okay it's Number 38 ! LOL, btw satellite.....ur old ,jk help?

OpenStudy (anonymous):

old? you have no idea!

OpenStudy (anonymous):

sorry it's side ways

myininaya (myininaya):

so we three 3 different linear equations

myininaya (myininaya):

so lets work on the first one

OpenStudy (anonymous):

i will do #38

OpenStudy (anonymous):

okay :D

myininaya (myininaya):

we see that (-2,4) and (-4,-2) lie on the first line

myininaya (myininaya):

we can find the slope and the y-intercept using \[m=\frac{y_1-y_2}{x_1-x_2}=\frac{4-(-2)}{-2-(-4)}=\frac{6}{2}=3\] \[y-y_1=m(x-x_1)\] \[y-4=3(x-(-2))\] \[y=3x+6+4=>y=3x+10\] so this line y=3x+10 is defined (-inf,2]

myininaya (myininaya):

now the second line is a constant which is y=4 it is defined on (-2,2)

OpenStudy (anonymous):

three lines. they change at -1 and 1 between -1 and 1 the function is constant. it is 2 before -1 it is a line with slope 3 (over 1, up 3) and since it goes up to (-1,2) the equation is \[y-2=3(x-1)\] or \[y=3x+5\] so we know it will look like \[f(x) = \left\{\begin{array}{rcc} 3x + 4 & \text{if} & x \leq -1 \\ 2& \text{if} & -1<x < 1\\ \text{something}& \text{if} & x>1 \end{array} \right. \]

myininaya (myininaya):

now last line

OpenStudy (anonymous):

opps sorry for the picture flip, but it sounds right

myininaya (myininaya):

we see (2,4) and (4,-2) is on the line \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{4-(-2)}{2-4}=\frac{6}{-2}=-3\] \[y-y_1=m(x-x_1)\] \[y-(-2)=-3(x-4)\] \[y+2=-3x+12\] \[y=-3x+10\] defined on [2,inf)

OpenStudy (anonymous):

so this line y=3x+10 is defined (-inf,2] <------ how do u know it is closed circle?

OpenStudy (anonymous):

they match up there so it makes no difference

myininaya (myininaya):

you can do it either way it doesn't matter

OpenStudy (anonymous):

OHHH

myininaya (myininaya):

you can include for the constant line instead

OpenStudy (anonymous):

still i am wrong!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} 3x + 10 & \text{if} & x \leq -2 \\ 4 & \text{if} & -2<x < 2\\ -3x+10& \text{if} & x\geq 2 \end{array} \right.\]

myininaya (myininaya):

lol satellite

OpenStudy (anonymous):

why do they have to count by 2's?

OpenStudy (anonymous):

lol satellite is to eager to help so he/she keeps messing up

OpenStudy (anonymous):

i am going to go delete all my wrong posts.

OpenStudy (anonymous):

my excuse it i had to read it sideways. that, and i am old...

OpenStudy (anonymous):

lol maybe you should retire

myininaya (myininaya):

satellite is awesome don't make him retire

OpenStudy (anonymous):

yeah it is time. after labor day for sure

myininaya (myininaya):

don't leave me satellite

OpenStudy (anonymous):

it's just an option, you really dont have to if you don't want to

OpenStudy (anonymous):

actually i just wanted to show off my piecewise function.

OpenStudy (anonymous):

LOL, i wish i was smart,so i would actually have the ability to show off

OpenStudy (anonymous):

BTW ty all